OpenStudy (anonymous):

The resultant of two forces acting at right angles is √10N when the same two forces act at 60° the resultant is √13N calculate the forces

OpenStudy (abhisar):

For the first case i.e 90° \(\sf a^2+b^2 +abcos90\)=10

OpenStudy (abhisar):

\(\Rightarrow a^2+b^2=10\)

OpenStudy (abhisar):

are the two forces equal ?

OpenStudy (anonymous):

No they are not equal

OpenStudy (abhisar):

i ma not sure about this one. I am getting ab =10/13

OpenStudy (abhisar):

I don't know how to proceed further

OpenStudy (anonymous):

Ok..ty bro for u effort....i find by tomorrow by asking my lecturer

OpenStudy (abhisar):

yeah...u do that...till m trying my best

OpenStudy (anonymous):

Here is the image frm the notes .plz make me understand

OpenStudy (abhisar):

oh lol ! Solved

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar For the first case i.e 90° \(\sf a^2+b^2 +abcos90\)=10 \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

These r the images frm the solution bt i m nt getting it

OpenStudy (abhisar):

Yes i was doing a calculation mistake

OpenStudy (abhisar):

Now let's see

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar Since cos(90) = 0 =>\(\sf a^2+b^2+0=10\)... =>\(\sf a^2+b^2=10\) ok so far ? \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

Wan thw image of the question?

OpenStudy (abhisar):

no no Question is clear to me....i got the answers...

OpenStudy (abhisar):

Now in second case \(\sf a^2+b^2+2abcos60=13\\Since~cos60=1/2\\=>a^2+b^2+ab=13\) got it ?

OpenStudy (abhisar):

Wait i'll upload a pic of solution

OpenStudy (anonymous):

Ok

OpenStudy (abhisar):

done..uploading

OpenStudy (anonymous):

Ty...nd the value of two forces r 3N nd 1N

OpenStudy (anonymous):

Thanks abhisar....

OpenStudy (abhisar):

There is one more page wait !

OpenStudy (abhisar):